OpenStudy (anonymous):

Can someone solve this for me? -2(x-7) + 4x = 6

OpenStudy (anonymous):

-2x -2*(-7) +4x = 6 -2x +14 +4x = 6 2x = -8 x = -4

OpenStudy (anonymous):

x = -4

OpenStudy (anonymous):

-2(x-7) = -2x + 14 (by distributive property) So you have: -2x + 14 + 4x = 2x + 14 (by adding like terms) 2x + 14 = 6 becomes 2x = -8 (by algebra) 2x = -8 (algebra) x = -4 (algebra) Hope this helps. If it does, please choose good answer. Visit http://www.tutorsean.net for free math videos.